OpenStudy (anonymous):

practic 5-4 5 7/4 = 6 ?/4= plzzzzz help :(

OpenStudy (anonymous):

@ade1010

OpenStudy (fibonaccichick666):

can you simplify 7/4?

OpenStudy (fibonaccichick666):

turn it into a mixed number I mean

OpenStudy (anonymous):

idk we are still learning

OpenStudy (fibonaccichick666):

ok so, what do you thin you can try?

OpenStudy (anonymous):

2 17/12=3 ?/12=

OpenStudy (anonymous):

4 3/10+ 5 2/5

OpenStudy (anonymous):

3 7/8 +2 1/2

OpenStudy (anonymous):

5 2/3 + 3 1/4

OpenStudy (anonymous):

6 3/4 + 2 1/2

OpenStudy (anonymous):

1 1/12 + 3 1/6

OpenStudy (jordanloveangel):

wait with all of thoes? @shira44

OpenStudy (anonymous):

9 2/5 + 10 3/10

OpenStudy (jordanloveangel):

kk i sec

OpenStudy (anonymous):

7 1/3 + 5 11/12

OpenStudy (anonymous):

11 7/10 + 4

OpenStudy (jordanloveangel):

7 1/3 + 5 11/12 = 13.25

OpenStudy (jordanloveangel):

11 7/10 + 4=15.7

OpenStudy (anonymous):

2 2/3 + 4 3/4 and thx

OpenStudy (anonymous):

7 3/4 + 2 7/8

OpenStudy (jordanloveangel):

2 2/3 + 4 3/4=7.41666666667

OpenStudy (jordanloveangel):

7 3/4 + 2 7/8=10.625

OpenStudy (anonymous):

4 1/2 + 3 5/6

OpenStudy (jordanloveangel):

4 1/2 + 3 5/6=8.33333333333

OpenStudy (anonymous):

7 2/3 + 1 5/6

OpenStudy (anonymous):

2 1/4 + 4 3/5

OpenStudy (fibonaccichick666):

http://openstudy.com/code-of-conduct please read

OpenStudy (jordanloveangel):

sorry @FibonacciChick666

OpenStudy (fibonaccichick666):

don't be sorry to me, but with a 65 smart score, you should know better

OpenStudy (anonymous):

5 3/8 + 7 1/4

OpenStudy (jordanloveangel):

okay i wont do it again

OpenStudy (fibonaccichick666):

thank you

OpenStudy (jordanloveangel):

np

OpenStudy (jordanloveangel):

she's doing

OpenStudy (jordanloveangel):

@shira44 u still need help

OpenStudy (anonymous):

u can go i will get hel p somewhere else

OpenStudy (jordanloveangel):

umm o...k....y........

OpenStudy (anonymous):

Jordan can u give me a medal for the help i gave u last time and fan 2 plzzz. :)

OpenStudy (jordanloveangel):

yah sure y not @ade1010

OpenStudy (anonymous):

Ok thx @jordanloveangel fan too ?